# Large Language Models and Operations Research: A Structured Survey

### v1

@@ Abstract
Operations research (OR) provides fundamental methodologies for complex system decision-making, with established applications in transportation, supply chain management, and production scheduling. Traditional approaches, which depend on expert-based modeling and manual parameter adjustment, often face challenges in handling large-scale, dynamic, and multi-constraint problems. Recently, large language models (LLMs) have shown potential to address these limitations through semantic understanding, structured generation, and reasoning control. LLMs can translate natural language descriptions into mathematical models or executable code, generate heuristics, evolve algorithms, and directly tackle optimization tasks. This paper surveys recent progress on the integration of LLMs into OR, organizing methods into three main directions: automatic modeling, auxiliary optimization, and direct solving. It further reviews evaluation benchmarks and domain-specific applications, and summarizes key open issues such as unstable semantic-to-structure mapping, fragmented research progress, limited generalization, and insufficient evaluation systems. Finally, the survey outlines possible research avenues for advancing the role of LLMs in OR.

### v3

# Large Language Models in Operations Research: Methods, Applications, and Challenges

@@ Abstract
Operations research (OR) is a core methodology that supports complex system decision-making, with broad applications in transportation, supply chain management, and production scheduling. However, traditional approaches that rely on expert-driven modeling and manual parameter tuning often struggle with large-scale, dynamic, and multi-constraint problems, limiting scalability and real-time applicability. Large language models (LLMs), with capabilities in semantic understanding, structured generation, and reasoning control, offers new opportunities to overcome these challenges. It can translate natural language problem descriptions into mathematical models or executable code, generate heuristics, evolve algorithms, and directly solve optimization tasks. This shifts the paradigm from human-driven processes to intelligent human–AI collaboration. This paper systematically reviews progress in applying LLMs to OR, categorizing existing methods into three pathways: automatic modeling, auxiliary optimization, and direct solving. It also examines evaluation benchmarks and domain-specific applications, and highlights key challenges, including unstable semantic-to-structure mapping, fragmented research, limited generalization and interpretability, insufficient evaluation systems, and barriers to industrial deployment. Finally, it outlines potential research directions. Overall, LLMs demonstrate strong potential to reshape the OR paradigm by enhancing interpretability, adaptability, and scalability, paving the way for next-generation intelligent optimization systems.